# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 4.3.3 Continual Instruction Tuning (CIT)
Instruction Tuning (IT) is a technique used to refine the instruction-following capabilities of LLMs [353]. While LLMs are typically pre-trained on extensive and diverse corpora, they may struggle with specific tasks despite their general knowledge. Numerous studies have shown that IT can notably improve LLMs’ ability to follow textual instructions for particular tasks [353, 301, 125, 249, 218], leveraging the pre-existing knowledge within LLMs to bridge the gap between general and task-specific performance [302]. Additionally, IT enhances the intuitive interaction between humans and LLMs, providing a more natural interface and aligning LLM outputs more closely with human expectations and preferences [184].
@@
Continual Instruction Tuning (CIT). When the instruction tuning data comes in as a stream, this series of the fine-tuning tasks need to address the challenge of catastrophic forgetting of the general knowledge. CT0 [255] represents the inaugural study on continual learning in fine-tuned LLMs, utilizing the replay method during the IT process on the base T0 model. This model successfully learns 8 new tasks while maintaining robust performance on previously learned tasks. Subsequent research has focused on enhancing the replay method used during CIT. For instance, [103] improve replay efficiency by computing Key-part Information Gain (KPIG) on masked parts to dynamically select replay data, addressing the “half-listening” issue in instruction following. Similarly, SSR [113] uses the LLM to generate synthetic instances for replay, achieving superior or comparable performance with greater data efficiency than traditional methods.
@@
Other approaches include combining multiple continual techniques during CIT. DynaInst [205] merges parameter regularization with dynamic replay, selectively storing and replaying instances and tasks to enhance outcomes. InstructionSpeak [337] employs negative training and replay instructions to improve both forward transfer and backward transfer. Additionally, some methods incorporate Parameter Efficient Tuning (PET). Orthogonal Low-Rank Adaptation (O-LoRA) learns new tasks within an orthogonal subspace while preserving LoRA parameters for previous tasks [291]. Shared Attention Framework (SAPT) combines a PET block with a selection module via a Shared Attentive Learning & Selection module, and tackles catastrophic forgetting and knowledge transfer concurrently [363]. While regularization-based and architectural-based methods require additional parameter storage and GPU memory during fine-tuning, which potentially leads to higher training costs, replay-based methods remain popular due to their simplicity and data efficiency [294]. These methods are particularly favored in traditional continual learning scenarios for tuning LLMs.
CIT vs Conventional CL. Both CIT and conventional CL aim to enable LLMs to acquire new tasks or information over time while retaining previously learned knowledge. However, CIT specifically utilizes rich natural language instructions to enhance LLMs’ ability to follow human instructions [358]. The natural-language encoding of task information enables the potential for positive forward transfer when semantically similar tasks are encountered in the data stream. This scenario is typically challenging to engineer manually in conventional continual learning setups. In contrast, conventional CL focuses on broader knowledge acquisition and is not limited to instruction-based learning. While CIT is predominantly applied within LLMs, conventional CL is employed across various fields, including vision, multimodal models and robotics. Regarding challenges, both approaches contend with catastrophic forgetting; however, CIT additionally concentrates on refining instruction-following capabilities to better interact with human needs. Conversely, conventional CL emphasizes building robust generalizability to manage variations within and across tasks [288].
4.3.4 Continual Model Refinement (CMR)
@@
Like humans, LLMs are prone to errors, such as inaccurate translations or outdated information [60]. To keep LLMs updated with the evolving factual knowledge, directly fine-tuning the model to correct these mistakes can be time-consuming and may disrupt its performance on previously learned tasks. To overcome these challenges, Model Refinement (also known as Model Editing) is proposed. This approach aims to rectify the model’s errors while preserving its performance on other inputs, with only moderate computing resources [264, 60, 203, 98, 117, 204, 96]. The concept of model editing was initially explored in [264], which introduced a “reliability-locality-efficiency” principle and proposed a gradient descent editor to address it efficiently. Subsequent research, such as [60] and [203], extended this principle to edit factual knowledge in BERT-based language models and larger models like GPT-J-6B [285] and T5-XXL [235], respectively, using gradient decomposition. These approaches typically update a subset of model parameters to alter the labels of specific inputs. Additionally, memory-based models, as discussed in [204] and [96], incorporate editing through retrieval mechanisms.
Model Refinement as a Special Form of Continual Learning. Although model refinement and continual learning are typically treated as separated fields, they share fundamental similarities. The principle of locality in model refinement, ensuring that new knowledge doesn’t disrupt responses to other inputs, aligns with the non-forgetting objective of continual learning. Furthermore, the principles of reliability—effectively updating the model—and efficiency are crucial in both model refinement and continual learning Thus, the problem of model refinement can be viewed as akin to continual learning, where a small batch of updated samples $\{({\bm{x}}_{e},\widehat{y}_{e})\}$ represents a new task. Alternatively, one can consider the stream of such samples for updating as a specialized form of online continual learning [335, 189, 228].
@@
Continual Model Refinement (CMR). Recent works have combined continual learning with model refinement, resulting in a new problem termed continual model refinement (CMR). This concept extends model refinement horizontally, presenting updated sample pairs ${({\bm{x}}_{e},y_{e},\widehat{y}_{e})}^{e=1}_{N}$ sequentially as a stream. [162] initially introduces this idea as continual model refinement, evaluating various continual learning methods with a dynamic sampling algorithm. To solve CMR problem, many methods employ a retrieval mechanism. For instance, [96] uses hidden activations of the language model as a “key” to activate updated parameters only when input $x_{0}$ resembles updated sample pairs; [339] improves this approach’s efficiency by integrating LoRA [109]; [58] augments the LLM with an external episodic memory, modeling CMR as an ongoing memory refresh. Meanwhile, some methods focus solely on updating a subset of model parameters. For example, [108] addresses the issue of “toxicity buildup and flash” in single-editing methods like ROME [196], adapting it to a continual context with a knowledge-aware layer selection algorithm.
@@
While all these works pioneer research in CMR, the exploration of continual model refinement for LLMs remains open. [97] highlighted a potential problem: the location for storing the fact may not coincide with the best place for editing it. This challenges the classical "locate and edit" paradigm used by several model editing methods [196, 197], and could become a significent concern for CME [108]. Other questions, including whether such problem setting fits LLMs and whether more memory/computationally efficient methods of CMR could be developed for LLMs, are yet to be explored and answered.
4.3.5 Continual Model Alignment (CMA)
@@
Model Alignment (MA) is a crucial concept in the development and deployment of AI systems, ensuring that their actions and outputs align with human values, ethics, and preferences. It can be defined as the process of adjusting the objectives and functioning of an AI system to achieve such goals, involving a combination of mathematical models, algorithmic adjustments, and iterative feedback to refine AI behavior [218, 234]. MA can be broadly categorized into two types: Reinforcement Learning-based (RL-based) and Supervised Learning-based (SL-based). In the RL-based approach, as discussed in [315, 218, 252], models are trained to make decisions reinforced by human feedback, using a reward system to guide them towards desirable outcomes. Conversely, the SL-based approach, as seen in [104, 234, 122], directly trains models on datasets of human preferences, aligning their output with demonstrated human values. Both approaches leverage a combination of algorithmic learning techniques and human feedback to progressively refine and align model behaviors.
When language models undergo the phase of model alignment (MA) to align with specific ethical or normative standards, the performance degradation occurs. In [165], the authors refer to this phenomenon of general knowledge forgetting induced by model alignment as the “Alignment Tax”. Notably, even a single stage of MA can diminish the model’s performance capabilities, as it restricts the model’s responses to a narrower subset of its training distribution. This paper focuses on understanding and reducing the impact of the Alignment Tax in the process of refining AI models, highlighting the balance between alignment and maintaining broad model capabilities.
Continual Model Alignment (CMA). Continual Model Alignment (CMA) aims to continuously refine LLMs to align with evolving human values, ethics, and data. The significance of CMA lies in its capacity to ensure that LLMs retain their relevance, accuracy, and ethical alignment over time. This ongoing adjustment is essential to navigate the complexities introduced by concept drift, the evolution of data, and shifts in societal values. The static nature of LLM training on historical data sets can lead to discrepancies between the models’ outputs and current factual accuracies, societal norms, and standards, making CMA a crucial process for maintaining their adaptability and alignment with contemporary contexts [275]. Despite its importance, CMA faces several challenges: (i) scalability and resource intensity issue: continually updating LLMs requires significant computational resources and human oversight, posing scalability issues for the practitioners; (ii) ensuring ethical alignment: balancing model updates to reflect societal changes without introducing or perpetuating biases remains a complex issue; (iii) data privacy and security: continuously integrating new data into LLMs raises concerns regarding data privacy, security, and the potential for misuse.
Emerging strategies in CMA emphasize the optimization of LLMs for enhanced adaptability, addressing the necessity for these models to accommodate continuous changes in language use, information validity, and societal expectations. Notably, innovative research efforts such as those detailed in [165], and [229], highlight the development of methodologies designed to minimize the challenges associated with the continuous realignment of LLMs. These studies underscore the importance of implementing specific optimization strategies to foster the adaptability of LLMs, ensuring their outputs remain ethically attuned and factually relevant in the face of dynamic societal and informational landscapes.
Likewise, there are two types of CMA frameworks: RL-based and SL-based. In the realm of RL-based CMA, two significant contributions have been noted: [165] explores methodologies for reducing the overhead associated with continual alignment through efficient reinforcement learning techniques, and [352], which suggests a framework for applying continual learning principles to reinforcement learning with human feedback, potentially mitigating the alignment tax over time. Employing reinforcement learning for CMA focus on developing efficient reinforcement learning techniques that integrate continual learning principles and human feedback to dynamically maintain alignment with evolving human values while minimizing computational overhead and mitigating forgetting. For SL-based CMA, [351] presents an innovative approach by integrating supervised continual learning techniques with the process of aligning AI systems through direct training on evolving datasets of human preferences. This methodology promises a more sustainable and adaptable model for maintaining alignment with human values over extended periods and across various contexts. The use of supervised continual learning suggests a focus on preventing catastrophic forgetting, a common challenge where a model loses its ability to perform previously learned tasks upon learning new ones.
In summary, both RL-based and SL-based frameworks need to address forgetting, especially in scenarios where the model continuously integrates new information. Effective strategies might include techniques (such as EWC [140], experience replay [38], and dynamic re-weighting [165]), which help the model retain old knowledge while integrating new insights. Future research in CMA aims to develop more efficient, automated processes for model updates, better mechanisms for ethical oversight, and innovative solutions to balance model relevance with privacy and security concerns. Existing streams of research highlight the importance of developing adaptable, efficient, and robust AI systems that can continually align with human values without substantial losses in performance or increased computational costs. Future research could explore hybrid models that combine RL-based and SL-based approaches, potentially offering a more holistic framework for continual model alignment.
4.3.6 Continual Multimodal Large Language Models (CMLLMs)
@@
Multi-modal LLMs have garnered significant attention for their capacity beyond single modality. These models integrate data from multiple modalities, like texts, images and videos to enhance real-world information comprehension [222, 148]. Typically, MLLMs consist of modality-specific sub-modules such as pre-trained vision encoders, large language models, and projectors for cross-model alignment. This alignment is essential for MLLMs to fuse the diverse data types and promote their comprehension. For example, MiniGPT-4 [370] utilizes a linear projector to align frozen vision encoders and language models; LLaVA [167] employs a simple linear layer to connect image features and instruction into the word embedding space. Currently, all existing MLLMs are pre-trained on large scale multi-modal datasets and then fine-tuned on specific small downstream datasets, which constitutes the training process [56, 154].
@@
Several existing studies have explored the causes of catastrophic forgetting when continually training MLLMs. [364] performs singular value decomposition on input embeddings, revealing a significant disparity among different input embeddings. This discrepancy causes the model to learn irrelevant information for previously trained tasks, resulting in catastrophic forgetting and negative forward transfer. [349] observes that minority collapse may lead to catastrophic forgetting, when the imbalance ratio between majority and minority classes approaches infinity during fine-tuning. It further identifies hallucination as a contributing factor to performance degradation in MLLMs.
Continual Pre-Training MLLMs. Training an MLLM to be updated with the changing world from scratch is resource-intensive, requiring considerable time and cost. While continual learning offers a promising solution to this problem, trivially applying past methods is not advisable, given the distinct structure of MLLMs that includes modality-specific sub-modules and cross-model alignment. Currently, there is an apparent lack of continual pre-training for MLLMs, and further exploration into continual and joint pre-training of MLLMs is necessary.
@@
Continual Fine-Tuning MLLMs. In contrast to traditional continual learning methods that involve full-model fine-tuning for new tasks, continual fine-tuning for MLLMs focuses on refining specific layers when adapting to new tasks [349, 100, 364, 41, 371]. Given the strong capabilities of pre-trained models, training specific layers suffices, and can simultaneously reduce computational demands. [362] additionally considers an continual learning scenario, Continual Missing Modality Learning (CMML), where different modalities are emerging throughout the incremental learning stages. All the aforementioned studies collectively indicate that MLLMs still suffer from catastrophic forgetting, which manifests in two ways: along the direction of vertical continuity, a performance decline on pre-trained tasks following fine-tuning for downstream tasks; and along the axis of horizontal continuity, a performance degrade on previously fine-tuned tasks after fine-tuning for new tasks. [364] also observes negative forward transfer, where the performance of unseen tasks degrades when learning new tasks, indicating a decline in model generalization capability.
@@
Continual Learning MLLMs. While traditional CL methods are applicable, some may not yield optimal results, as evidenced by various experiments [100, 364]. For instance, [100] observes a consistent efficacy of replay-based and model expansion strategies across diverse scenarios of continual fine-tuning MLLMs, but regularization-based methods only perform well on models that have been jointly instruction-tuned on multiple tasks. Other works seek to develop ad-hoc solutions for continual learning MLLMs. [100] proposes EProj to expand the projection layer in MLLMs for each new task and utilizes task-similarity-informed regularization (TIR) to enhance performance. [364] introduces Fwd-Prompt, a prompt tuning method that projects prompt gradient to both the residual space and the pre-trained subspace to minimize the interference between tasks and reuse pre-trained knowledge respectively, fostering positive forward transfer without relying on previous samples. [371] focuses on the forgetting of the pre-trained MLLMs after fine-tuned on specific tasks and proposes model tailor to compensate the selected subset that are critical for enhancing target task performance. [362] presents a novel method named Reconstruct before Query (RebQ), leveraging the multi-modal knowledge from a pre-trained model to reconstruct the absent information for the missing modality. Recently, MoE (Mixture-of-Experts) framework has gained attention which resembles the architecture-based methods in CL. It provides the model with the ability to learn different intentions from distinct experts, e.g., [41] first introduces MoELoRA to fine-tune LLaVA, effectively mitigate the catastrophic forgetting of MLLMs in CoIN and the results demonstrate the effectiveness.
In concluding remarks on Continual Learning for MLLMs, the role of templates in instruction tuning emerges as crucial. As highlighted by [41], employing similar templates across tasks proves more advantageous, aiding in knowledge retention and forgetting mitigation. This approach fosters task-specific learning, reducing reliance on common knowledge prone to forgetting in sequential contexts. In addition, it is noteworthy that the forgetting induced by the gap between tasks is more critical than the forgetting induced by the distributional gap between datasets. According to [349], moderate fine-tuning is advantageous for non-fine-tuned tasks, excessive fine-tuning ultimately leads to catastrophic forgetting in these tasks. [100] discovers the multi-task joint instruction tuning at the beginning state can facilitate the model’s continual learning ability and mitigate forgetting. Overall, continual learning in MLLMs holds promise, but further research is needed to fully realize its potential.
5 Evaluation Protocols and Datasets

### v2

@@ 4.3.3. Continual Instruction Tuning (CIT)
While LLMs are typically pre-trained on extensive and diverse corpora, they may struggle with specific tasks such as instruction following despite their general knowledge. Numerous studies have shown that Instruction Tuning (IT) can notably improve LLMs’ ability to follow textual instructions (zhang2024instruction, ; wei2021finetuned, ; jiang2024instructiontuned, ; sanh2022multitask, ; ouyang2022rlhf, ), leveraging the pre-existing knowledge within LLMs to bridge the gap between general and task-specific performance (wei2022finetuned, ). Recent works like WizardLM (xu2023wizardlm, ) and CodecLM (wang2024codeclm, ) further tailor synthetic data to steer LLMs’ behavior through IT. Additionally, IT enhances the interaction between humans and LLMs, providing a more natural interface and aligning LLM outputs more closely with human expectations and preferences (luo2023empirical, ).
@@
When the instruction tuning data comes in as a stream, forgetting of the previously learned instructions should be addressed. CT0 (scialom2022fine, ) represents the inaugural study on Continual Instruction Tuning (CIT) of LLMs, applying the replay method on the base T0 model throughout the process. Many subsequent studies focus on enhancing the replay method used during CIT. For instance, (he2024dont, ) improve replay efficiency by computing Key-part Information Gain (KPIG) on masked parts to dynamically select replay data, addressing the “half-listening” issue in instruction following. Similarly, SSR (huang2024mitigating, ) uses the LLM to generate synthetic instances for replay, achieving superior or comparable performance to traditional methods at a lower cost.
@@
Other approaches introduce multiple CL techniques during CIT. DynaInst (mok2023large, ) merges parameter regularization with dynamic replay, selectively storing and replaying instances and tasks to enhance outcomes. InstructionSpeak (yin2022contintin, ) employs negative training and replay instructions to improve both forward transfer and backward transfer. Some methods incorporate PEFT. Orthogonal Low-Rank Adaptation (O-LoRA) learns new tasks within an orthogonal subspace while preserving LoRA parameters for previous tasks (wang2023orthogonal, ) to minimize the interference among different tasks. Shared Attention Framework (SAPT) combines a PET block with a selection module via a Shared Attentive Learning & Selection module, tackling catastrophic forgetting and knowledge transfer concurrently (zhao2024sapt, ). While regularization-based and architectural-based methods require additional parameter storage and GPU memory, together with replay-based methods they remain for CIT due to the simplicity and effectiveness (wang2024inscl, ).
4.3.4. Continual Model Refinement (CMR)
@@
Like humans, LLMs are prone to errors, such as inaccurate translations or outdated information (de2021editing, ). Directly fine-tuning the model to correct these mistakes may disrupt its performance on previously learned tasks. To overcome these challenges, Model Refinement (also known as Model Editing) is proposed, which aims to rectify the model’s errors while preserving its performance on other inputs, with only moderate computing resources (sinitsin2020editable, ; de2021editing, ; fast_edit, ; hase2021language, ; huang2023transformer, ; mitchell2022memory, ; hartvigsen2023aging, ). The concept of model editing was initially explored in (sinitsin2020editable, ), which introduced a “reliability-locality-efficiency” principle and proposed a gradient descent editor to address it efficiently. Subsequent research, such as (de2021editing, ) and (fast_edit, ), extended this principle to edit factual knowledge in BERT-based language models and larger models like GPT-J-6B (gpt-j, ) and T5-XXL (raffel2020exploring, ), respectively, using gradient decomposition. These approaches typically update a subset of model parameters to alter the labels of specific inputs. Additionally, memory-based models, as discussed in (mitchell2022memory, ) and (hartvigsen2023aging, ), incorporate editing through retrieval mechanisms.
@@
The concept of Continual Model Refinement (CMR) extends model refinement horizontally, presenting updated sample pairs ${({\bm{x}}_{e},y_{e},\widehat{y}_{e})}^{e=1}_{N}$ sequentially as a stream. (lin2022continual, ) initially introduces this idea, evaluating various CL methods with a dynamic sampling algorithm. Many CMR methods employ a retrieval mechanism. For instance, (hartvigsen2023aging, ) uses hidden activations of the language model as a “key” to activate updated parameters only when input $x_{0}$ resembles updated sample pairs; (yu2023melo, ) improves this approach’s efficiency by integrating LoRA (hu2021lora, ); (das2024larimar, ) augments the LLM with an external episodic memory, modeling CMR as an ongoing memory refresh. Meanwhile, some methods focus solely on updating a subset of model parameters. For example, (hu2024wilke, ) addresses the issue of “toxicity buildup and flash” in single-editing methods like ROME (meng2022locating, ), adapting it to the CL context with a knowledge-aware layer selection algorithm. WISE (wang2024wise, ) addresses the “impossible triangle” of reliability, locality, and generalization in existing lifelong model refinement methods. It introduces a side memory system that enables knowledge sharding and merging, successfully achieving all three objectives simultaneously.
@@
While all these works pioneer research in CMR, the exploration of CMR of LLMs remains open. (hase2023does, ) highlights a potential problem: the location for storing the fact may not coincide with the best place for editing it. This challenges the classical “locate and edit” paradigm used by several existing methods (meng2022locating, ; meng2022mass, ), and could become a significent concern for CMR (hu2024wilke, ). Other questions, including whether such problem setting fits LLMs and whether more memory/computationally efficient methods of CMR could be developed for LLMs, are yet to be answered.
4.3.5. Continual Model Alignment (CMA)
@@
Model Alignment (MA) ensures AI systems’ actions and outputs align with human values, ethics, and preferences (ouyang2022rlhf, ; rafailov2024dpo, ). MA can be broadly categorized into two types: Reinforcement Learning-based (RL-based) and Supervised Learning-based (SL-based). RL-based approaches (wu2022survey, ; ouyang2022rlhf, ; schulman2017proximal, ) are trained to make decisions reinforced by human feedback, using a reward system to guide them towards desirable outcomes. Conversely, the SL-based approaches (hendrycks2023aligning, ; rafailov2024dpo, ; ji2024ai, ) directly train models on datasets of human preferences, aligning their output with demonstrated human values. Both approaches leverage a combination of algorithmic learning techniques and human feedback to progressively refine the model behavior. When LLMs undergo the phase of MA, vertical forgetting of previous knowledge usually occurs. In (lin2024mitigating, ), the authors refer to this phenomenon of catastrophic forgetting induced caused by MA as the “Alignment Tax”. Notably, even a single stage of MA can diminish the model’s performance capabilities, as it restricts the model’s responses to a narrower subset of the training distribution.
Continual Model Alignment (CMA) aims to continuously refine LLMs to align with evolving human values, ethics, and data. The static nature of LLM training on historical data sets can lead to discrepancies between the models’ outputs and current factual accuracies, societal norms, and standards, making CMA a crucial process for maintaining their adaptability and alignment with contemporary contexts (taori2023alpaca, ). Likewise, there are two types of CMA frameworks: RL-based and SL-based. In the realm of RL-based CMA, two significant contributions have been noted. (lin2024mitigating, ) identifies the conflicts between the existing CL techniques and RLHF, and proposes Adaptive Model Averaging (AMA), adaptively finding appropriate ratios for the combination of model layers to gain maximal rewards with minimal tax; Continual Proximal Policy Optimization (CPPO) (zhangcppo, ) proposes a weighting strategy for different examples deciding its usage of policy enhancement or knowledge retention, mitigating the alignment tax over time. For SL-based CMA, Continual Optimal Policy Fitting (COPF) (zhang2023copf, ) presents a solution adapted from the Direct Policy Optimization (DPO) (rafailov2024direct, ), solving its potential risks of sub-optimal policy fitting and over-optimization in the context of CMA.
4.3.6. Continual Multimodal Large Language Models (CMLLMs)
@@
Multi-modal LLMs integrate data of multiple modalities, like texts, images and videos to enhance real-world information comprehension (peng2023kosmos2, ; li2023otter, ). Typically, MLLMs consist of modality-specific sub-modules such as pre-trained vision encoders, large language models, and projectors for cross-model alignment. This alignment is essential for MLLMs to fuse the diverse data types and promote their comprehension.
@@
Continually training multi-modal models like CLIP (radford2021learning, ) has been long studied (zheng2023preventing, ; ni2023continual, ; wu2024building, ; khan2021personalizing, ; liu2023class, ; garg2024tic, ; li2024coleclip, ; jha2024clap4clip, ; yu2024select, ; yu2024boosting, ), while the problem of continually training MLLMs still remains underexplored. Several existing studies have investigated the causes of catastrophic forgetting when continually training MLLMs. (zheng2024antiforgetting, ) performs singular value decomposition on input embeddings, revealing a significant disparity among different input embeddings. This discrepancy causes the model to learn irrelevant information for previously trained tasks, resulting in catastrophic forgetting and negative forward transfer. (zhai2023investigating, ) observes that minority collapse may lead to catastrophic forgetting, when the imbalance ratio between majority and minority classes approaches infinity during fine-tuning. It further identifies hallucination as a contributing factor to performance degradation in MLLMs.
@@
Continual Fine-Tuning MLLMs. In contrast to traditional continual learning methods that involve full-model fine-tuning for new tasks, continual fine-tuning for MLLMs focuses on refining specific layers when adapting to new tasks (zhai2023investigating, ; he2023continual, ; zheng2024antiforgetting, ; chen2024coin, ; zhu2024model, ). Given the strong capabilities of pre-trained models, training specific layers suffices, and can simultaneously reduce computational demands. (zhao2024reconstruct, ) additionally considers an continual learning scenario, Continual Missing Modality Learning (CMML), where different modalities are emerging throughout the incremental learning stages. All the aforementioned studies collectively indicate that MLLMs still suffer from catastrophic forgetting, which manifests in two ways: along the direction of vertical continuity, a performance decline on pre-trained tasks following fine-tuning for downstream tasks; and along the axis of horizontal continuity, a performance degrade on previously fine-tuned tasks after fine-tuning for new tasks. (zheng2024antiforgetting, ) also observes negative forward transfer, where the performance of unseen tasks degrades when learning new tasks, indicating a decline in model generalization capability.
@@
While traditional CL methods are applicable, some may not yield optimal results, as evidenced by various experiments (he2023continual, ; zheng2024antiforgetting, ). For instance, (he2023continual, ) observes a consistent efficacy of replay-based and model expansion strategies across diverse scenarios of continual fine-tuning MLLMs, but regularization-based methods only perform well on models that have been jointly instruction-tuned on multiple tasks. Other works seek to develop ad-hoc solutions for continual learning MLLMs. (he2023continual, ) proposes EProj to expand the projection layer in MLLMs for each new task and utilizes task-similarity-informed regularization (TIR) to enhance performance. (zheng2024antiforgetting, ) introduces Fwd-Prompt, a prompt tuning method that projects prompt gradient to both the residual space and the pre-trained subspace to minimize the interference between tasks and reuse pre-trained knowledge respectively, fostering positive forward transfer without relying on previous samples. (zhu2024model, ) focuses on the forgetting of the pre-trained MLLMs after fine-tuned on specific tasks and proposes model tailor to compensate the selected subset that are critical for enhancing target task performance. (zhao2024reconstruct, ) presents a novel method named Reconstruct before Query (RebQ), leveraging the multi-modal knowledge from a pre-trained model to reconstruct the absent information for the missing modality. Recently, MoE (Mixture-of-Experts) framework has gained attention which resembles the architecture-based methods in CL. It provides the model with the ability to learn different intentions from distinct experts, e.g., (chen2024coin, ) first introduces MoELoRA to fine-tune LLaVA, effectively mitigate the catastrophic forgetting of MLLMs in CoIN and the results demonstrate the effectiveness.
5. Evaluation Protocols and Datasets